OpenStudy (anonymous):

If f(x) = x + 5, what does f(8) represent? A. The value of x when (x + 5) = 8 B. The value of 8f(x) when x = 8 C. The value of f(x) when (x + 5) = 8 D. The value of (x + 5) when x = 8

OpenStudy (anonymous):

@KamiBug

OpenStudy (anonymous):

D because it's telling you to plug in 8 to the problem

OpenStudy (ganpat):

what do u think ?

OpenStudy (dmndlife24):

It is the value of (x + 5) when x = 8

OpenStudy (anonymous):

value of (x + 5) when x = 8

OpenStudy (uri):

basically x =8 like f(x) =x + 5 f(8) =8+5

OpenStudy (anonymous):

so D ?

OpenStudy (dmndlife24):

8 represents what you should substitute x for... f(8) = 8 + 5 f(8) = 13 So yes...I would say D.

OpenStudy (uri):

Yeah D, @IndianBoyNiko

OpenStudy (anonymous):

A sequence has its first term equal to 4, and each term of the sequence is obtained by adding 2 to the previous term. If f(n) represents the nth term of the sequence, which of the following recursive functions best defines this sequence? A. f(1) = 2 and f(n) = f(n - 1) + 4; n > 1 B. f(1) = 4 and f(n) = f(n - 1) + 2n; n > 1 C. f(1) = 2 and f(n) = f(n - 1) + 4n; n > 1 D. f(1) = 4 and f(n) = f(n - 1) + 2; n > 1

OpenStudy (anonymous):

@uri @dmndlife24